OpenStudy (anonymous):

f(x) = 2x + 6, g(x) = 4x2 Find (f + g)(x).

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

(f+g)(x) is the same as f(x) + g(x)

jimthompson5910 (jim_thompson5910):

similar example: if f(x) = 2x and g(x) = x^2 - 3, then, (f+g)(x) = f(x) + g(x) (f+g)(x) = 2x + x^2 - 3 (f+g)(x) = x^2 + 2x - 3

OpenStudy (anonymous):

so it would be 4x^2+2x+6? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

wow thanks (: @jim_thompson5910

jimthompson5910 (jim_thompson5910):

np